# Onboard MuJoCo-based Model Predictive Control for Shipboard Crane with Double-Pendulum Sway Suppression

### 17 Mar 2026 v1

@@ II-B Sampling-Based Model Predictive Control
Our implementation utilizes the CEM planner provided in the MJPC framework to manage these rollouts asynchronously, as outlined in Algorithm 1. At each time $t$, the planner receives a state estimate $\hat{\mathbf{x}}(t)$ and samples a batch of $N$ control trajectories from a parametric distribution $\pi_{\theta}(\mathbf{U})$ - typically a diagonal Gaussian $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. These trajectories are represented as splines with $K$ knots, which serve as the decision variables for the optimization.
@@ IV-B1 Cross-Entropy Method for Real-Time Control
Trajectory sample size ($N$) 20
@@ VI-A Payload Sway Damping and Target Tracking under Perturbation
To evaluate the performance of our MJPC controller, we design a target switching task using two targets mounted $1$ m apart on the motion platform, shown in Fig. 1. Starting over one of the targets, we allow $20$ seconds for the controller to reach its new target and damp payload sway, after which we switch targets. We collect data from $10$ trajectories between the two targets with a total duration of $200$ seconds.
In Fig. 3, we display aggregated trajectories over all replications of the payload’s $xy$ position error relative to the target (top) and the tilt angle between the payload longitudinal axis and the z-axis of the global frame (bottom). To summarize the controller’s ability to maintain payload stability over the target, Table V aggregates the mean $xy$ position error over the $10$ to $20$-second interval, once the payload has reached the target.
@@ VI-B Feasibility of Real-Time MJPC with CEM Planning
We first analyze convergence of the planner in simulation. We configure MJPC to generate plans using a varying number of CEM iterations per planning cycle, ranging from $1$ to $100$. For each configuration, average the nominal return (lowest trajectory cost prior to the action update) over $50$ independent runs. As shown in Fig. 4, CEM demonstrates clear convergence, with the nominal cost decreasing sharply and eventually plateauing after approximately $60$ iterations. However, each additional iteration linearly increases planning time (dashed black line in Fig. 4, and exceeding $10$ iterations violates the $20$ Hz real-time control frequency constraint.
To establish practical feasibility, we replicate this analysis on the physical hardware, restricting the range from $1$ to $10$ iterations. For each configuration, we deploy the controller to maintain the payload over a stationary target and record the $xy$ position error over $1$ minute. Fig. 5 shows the mean position error for slow, medium and fast platform speeds. Too few iterations produce under-converged policies with high tracking error, while too many introduce latency that destabilises the system. Operating at five planning iterations strikes the balance: the resulting replanning rate (${\sim}40$ Hz) provides sufficient policy refinement to suppress double-pendulum sway while remaining well above the crane control frequency.
@@ VII Conclusion and Future Work
Future work will address current limitations regarding disturbance prediction and sensing. To improve responsiveness to sudden aperiodic disturbances, we plan to replace the current periodic history assumption with a state estimator, such as a Kalman filter. Furthermore, to eliminate reliance on external motion capture, we will implement onboard sensor fusion using IMUs, RGB-D cameras, and LiDAR to enhance reliability in outdoor settings. Finally, we aim to extend the MuJoCo model with an additional hinge joint to capture and suppress the rotational spin of the payload about the cable axis, a dynamic currently unmodelled in the $xy$ plane representation.

### 25 Mar 2026 v2

@@ II-B Sampling-Based Model Predictive Control
Our implementation utilizes the CEM planner provided in the MJPC framework to manage these rollouts asynchronously, as outlined in Algorithm 1. At each time $t$, the planner receives a state estimate $\hat{\mathbf{x}}(t)$ and samples a batch of NN control trajectories from a parametric distribution $\pi_{\theta}(\mathbf{U})$ - typically a diagonal Gaussian $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. These trajectories are represented as splines with $K$ knots, which serve as the decision variables for the optimization.
@@ IV-B1 Cross-Entropy Method for Real-Time Control
Trajectory sample size (NN) 20
@@ VI-A Payload Sway Damping and Target Tracking under Perturbation
To evaluate the performance of our MJPC controller, we design a target switching task using two targets mounted 11 m apart on the motion platform, shown in Fig. 1. Starting over one of the targets, we allow 2020 seconds for the controller to reach its new target and damp payload sway, after which we switch targets. We collect data from $10$ trajectories between the two targets with a total duration of $200$ seconds.
In Fig. 3, we display aggregated trajectories over all replications of the payload’s xyxy position error relative to the target (top) and the tilt angle between the payload longitudinal axis and the z-axis of the global frame (bottom). To summarize the controller’s ability to maintain payload stability over the target, Table V aggregates the mean xyxy position error over the $10$ to 2020-second interval, once the payload has reached the target.
@@ VI-B Feasibility of Real-Time MJPC with CEM Planning
We first analyze convergence of the planner in simulation. We configure MJPC to generate plans using a varying number of CEM iterations per planning cycle, ranging from 11 to $100$. For each configuration, average the nominal return (lowest trajectory cost prior to the action update) over $50$ independent runs. As shown in Fig. 4, CEM demonstrates clear convergence, with the nominal cost decreasing sharply and eventually plateauing after approximately $60$ iterations. However, each additional iteration linearly increases planning time (dashed black line in Fig. 4, and exceeding $10$ iterations violates the 2020 Hz real-time control frequency constraint.
To establish practical feasibility, we replicate this analysis on the physical hardware, restricting the range from 11 to $10$ iterations. For each configuration, we deploy the controller to maintain the payload over a stationary target and record the xyxy position error over 11 minute. Fig. 5 shows the mean position error for slow, medium and fast platform speeds. Too few iterations produce under-converged policies with high tracking error, while too many introduce latency that destabilises the system. Operating at five planning iterations strikes the balance: the resulting replanning rate (${\sim}40$ Hz) provides sufficient policy refinement to suppress double-pendulum sway while remaining well above the crane control frequency.
@@ VII Conclusion and Future Work
Future work will address current limitations regarding disturbance prediction and sensing. To improve responsiveness to sudden aperiodic disturbances, we plan to replace the current periodic history assumption with a state estimator, such as a Kalman filter. Furthermore, to eliminate reliance on external motion capture, we will implement onboard sensor fusion using IMUs, RGB-D cameras, and LiDAR to enhance reliability in outdoor settings. Finally, we aim to extend the MuJoCo model with an additional hinge joint to capture and suppress the rotational spin of the payload about the cable axis, a dynamic currently unmodelled in the xyxy plane representation.